OpenStudy (anonymous):

7x __ = 21 8 A.24 B.18.375 C.1.857 D.19

OpenStudy (anonymous):

multiply both sides by 8 and divide them by 7

OpenStudy (anonymous):

so do 8*8 and 7*8

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and then divide both bye seven

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no idk

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

7x is over a line and then 8 is on the bottom and 21 is on the side bye an equal sign

OpenStudy (solomonzelman):

\[\huge\color{blue}{\huge {\bbox[5pt,cyan,border:2px solid purple]{\frac{7x}{8}=21}}}\]multiply btoh sides times 8, and divide both sides by 7

OpenStudy (anonymous):

so 56/7 and 8/64 and i vget 8 for both

OpenStudy (solomonzelman):

no.

OpenStudy (solomonzelman):

\[24\]

OpenStudy (anonymous):

really wow

OpenStudy (solomonzelman):

\(\huge\color{ magenta}{\huge {\bbox[5pt,cyan,border:2px solid blue ]{Yup,~~it's~~24.}}}\)

OpenStudy (anonymous):

-6q+5.8=23.8 q=3 q=-3 q=1.83 q=4.93

OpenStudy (anonymous):

its B, I just worked it out

OpenStudy (anonymous):

@katalina13 anything else?

OpenStudy (anonymous):

which equation matches the statement "Th product of 3 and the difference of a number and 10 is 15 A.3+(x-10=15 B.3x-10=15 C.3(x-10=15 D.3x -- 15 10

OpenStudy (anonymous):

Well, what do you think it is

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

i have two more after this

OpenStudy (anonymous):

okay, Well, for this one, the product is the....

OpenStudy (anonymous):

the product is x right

OpenStudy (anonymous):

Do you know PEMDAS?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, thats good, that'll help you, Imma just tell you the answer cause you're busy, the answer is C

OpenStudy (anonymous):

thnx do you still have time for 2 more

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok i'll post them both the you can answer

OpenStudy (anonymous):

okay?lol

OpenStudy (anonymous):

ernsto loses 14.5 pounds in 7 months. now he ways 156.5 lb. which equation can be used to find how much he weighed before x+14.5=156.5 x-14.5=156.5 7x+14.5=156.5 7x-14.5=156.5 shannon opens a savings account with $50. strating with the frist week each week she deposits $8 into her account how many weeks does it take her to save $450

OpenStudy (anonymous):

56.25 weeks 62.5 weeks 52 weeks 50 weeks

OpenStudy (anonymous):

1. B 2. D

OpenStudy (anonymous):

Im pretty sure

OpenStudy (anonymous):

ok thanks so much your a life saver

OpenStudy (anonymous):

What did you have for the last two

OpenStudy (anonymous):

the same

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Well, I hope they were right, let me know please

OpenStudy (anonymous):

i will in private message

OpenStudy (anonymous):

Okay